# The new observations about the parameter-dependent Schrödinger-Poisson system

### v1

@@ 4.1. Cerami condition
$\begin{split}c+o_{n}(1)&=I_{\epsilon}(u_{n})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x\\ &\quad\quad-\frac{1}{3\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\left(1+\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(\delta|u_{n}|^{2}+\epsilon C_{\delta}|u_{n}|^{3}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\left(\frac{1}{2}-\frac{1}{3}\right)-\frac{1}{4}\left(\frac{1}{3}-\frac{1}{4}\right)-\delta\left(1+\frac{1}{3}\right)\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x\\ &\quad\quad+\left(\frac{1}{3}-\frac{1}{4}-C_{\delta}\epsilon\right)\int_{\mathbb{R}^{3}}|u_{n}|^{3}\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x.\end{split}$
Take $\delta=\frac{1}{64}$. Then, there exists a constant $\epsilon_{*}>0$ small enough such that when $0<\epsilon\leq\epsilon_{*}$, one has
$\left(\frac{1}{3}-\frac{1}{4}-C_{\frac{1}{64}}\epsilon\right)\geq 0$
@@
$c+o_{n}(1)\geq\frac{1}{8}\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x.$
@@
$\left(\frac{1}{2}-\frac{1}{3}\right)m\|v_{n}\|_{L^{2}(\mathbb{R}^{3})}^{2}\geq\frac{1}{16}.$
@@
On the other hand, it follows from assumptions $({\bf f})$ and $({\bf f^{\prime}})$ that
(4.2) $\tilde{F}(t)\geq 0\ ~\text{for any}~\epsilon\in(0,\frac{\delta_{0}^{2}}{4}).$
From (4.2) and the Fatou lemma, we have
@@
$\begin{split}&\indent\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\tilde{F}(\epsilon u_{n})+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}\text{d}x\\ &=\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4u_{n}^{2}}+\frac{\tilde{F}(\epsilon u_{n})}{\epsilon^{2}u_{n}^{2}}+\frac{mu_{n}^{2}}{2u_{n}^{2}}\right)v^{2}_{n}\text{d}x\\ &\geq\frac{1}{4}\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\to+\infty\end{split}$
as $n\to\infty$. Thus
@@
$\begin{split}c-1&\leq I_{\epsilon}(u_{n})\\ &=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x\\ &\leq\rho_{n}^{2}\left(\frac{1}{2}-\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\tilde{F}(\epsilon u_{n})+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}(x)\text{d}x\right)\to$-$\infty\end{split}$
as $n\to\infty$. This is impossible. Therefore, our claim is true.
From the boundedness of $\{u_{n}\}$ and $H^{1}_{V}(\mathbb{R}^{3})$ is compactly embedded in $L^{p}(\mathbb{R}^{3})(p\in[2,6))$, up to a sub-sequence we may assume
$\begin{split}u_{n}&\to u~\text{weakly in}~H^{1}_{V}(\mathbb{R}^{3});\\ u_{n}&\to u~\text{strongly in}~L^{p}(\mathbb{R}^{3})(p\in[2,6));\\ u_{n}&\to u~\text{a.e. in}~\mathbb{R}^{3}.\end{split}$
Next let $\epsilon_{**}=\min\{\epsilon_{*},\frac{\delta^{2}_{0}}{4}\}$, we can follow the same argument as Lemma 2 to show that $\{u_{n}\}$ has a convergent subsequence. ∎

### v3

@@ 4.1. Cerami condition
$\begin{split}c+o_{n}(1)&=I_{\epsilon,R}(u_{n})-\frac{1}{3}\langle I_{\epsilon,R}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\quad\quad-\frac{2}{3\epsilon^{2}}\eta_{R}^{\prime}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2})\text{d}x\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\text{d}x\\ &\quad\quad-\frac{1}{3\epsilon}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left(f\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\left(\eta_{R}\left(\left|u_{n}\right|\right)u_{n}+\eta_{R}^{\prime}\left(\left|u_{n}\right|\right)u_{n}^{2}\right)\right)\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\frac{4}{3}\left(\frac{1}{R}+C_{R}\epsilon^{p-2}\right)-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &=\frac{1}{6}\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\frac{1}{6}\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x-\frac{4}{3}\left(\frac{1}{R}+C_{R}\epsilon^{p-2}\right).\end{split}$
Then, there exists a constant $\epsilon^{\prime}(R)>0$ small enough such that when $0<\epsilon\leq\epsilon^{\prime}(R)$, one has
$\frac{1}{R}+C_{R}\epsilon^{p-2}\leq 1$
@@
$c+o_{n}(1)\geq\frac{1}{6}\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{1}{6}\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x.$
@@
$\|v_{n}\|_{L^{2}(\mathbb{R}^{3})}^{2}\geq\frac{1}{2m}.$
@@
On the other hand, it follows from assumptions $({\bf f})$ and $({\bf f^{\prime}})$ that for $\epsilon\in(0,\frac{\delta_{0}}{R})$,
(4.1) $F\left(\epsilon\eta_{R}\left(\left|t\right|\right)t\right)\geq 0\ ~\text{for all}~t\in\mathbb{R}.$
From (4.1) and the Fatou lemma, for $\epsilon\in(0,\frac{\delta_{0}}{R})$ we have
@@
$\begin{split}&\indent\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}\text{d}x\\ &\geq\int_{\mathbb{R}^{3}}\frac{\phi_{u_{n}}u^{2}_{n}}{4u_{n}^{2}}v^{2}_{n}\text{d}x\\ &\geq\frac{1}{4}\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\to+\infty,~~\text{as}~n\to\infty.\end{split}$
Thus
@@
$\begin{split}c-1&\leq I_{\epsilon,R}(u_{n})\\ &=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x\\ &=\rho_{n}^{2}\left(\frac{1}{2}-\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}\text{d}x\right)\\ &\to-\infty,~~\text{as}~n\to\infty.\end{split}$
This is impossible. Therefore, our claim is true.
From the boundedness of $\{u_{n}\}$ and $H^{1}_{V}(\mathbb{R}^{3})$ is compactly embedded in $L^{q}(\mathbb{R}^{3})(q\in[2,6))$, up to a sub-sequence we may assume
$\begin{split}u_{n}&\to u_{0}~\text{weakly in}~H^{1}_{V}(\mathbb{R}^{3});\\ u_{n}&\to u_{0}~\text{strongly in}~L^{q}(\mathbb{R}^{3})(q\in[2,6));\\ u_{n}&\to u_{0}~\text{a.e. in}~\mathbb{R}^{3}.\end{split}$
Next we prove the strong convergence of $u_{n}$ in $H_{V}^{1}\left(\mathbb{R}^{3}\right)$. To simplify the notation, let $\Psi_{\epsilon,R}\left(u\right):=\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x$.
Claim:
$\langle I^{\prime}_{\epsilon,R}(u_{n})-I^{\prime}_{\epsilon,R}(u_{0}),u_{n}-u_{0}\rangle$
$=\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}-\int_{\mathbb{R}^{3}}(\phi_{u_{n}}u_{n}-\phi_{u_{0}}u_{0})(u_{n}-u_{0})\text{d}x-m\int_{\mathbb{R}^{3}}(u_{n}-u_{0})^{2}\text{d}x$
$\quad-\langle\Psi^{\prime}_{\epsilon,R}(u_{n})-\Psi^{\prime}_{\epsilon,R}(u_{0}),u_{n}-u_{0}\rangle$
$\geq\frac{1}{2}\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}+o_{n}(1).$
Indeed, by the boundedness of ${u_{n}}$ in $H^{1}_{V}(\mathbb{R}^{3})$, we have
$\left|\int_{\mathbb{R}^{3}}\left(\phi_{u_{n}}u_{n}-\phi_{u_{0}}u_{0}\right)\left(u_{n}-u_{0}\right)\text{d}x\right|\leq o_{n}\left(1\right)\left\|u_{n}-u_{0}\right\|_{H^{1}_{V}(\mathbb{R}^{3})}=o_{n}\left(1\right)$
and
$\int_{\mathbb{R}^{3}}(u_{n}-u_{0})^{2}\text{d}x=o_{n}(1).$
Thus, we only need to prove that
$\langle\Psi_{\epsilon,R}^{\prime}(u_{n})-\Psi_{\epsilon,R}^{\prime}(u_{0}),u_{n}-u_{0}\rangle\leq\frac{1}{2}\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}+o_{n}(1).$
A direct computation shows that
$\displaystyle\langle\Psi^{\prime}_{\epsilon,R}(u),w\rangle=$ $\displaystyle\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}(\nabla u\nabla w+\tilde{V}(x)uw)\text{d}x\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u|)u)\text{d}x$
$\displaystyle+\frac{1}{\epsilon}\eta_{R}(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}f(\epsilon\eta_{R}(|u|)u)\left(\eta_{R}(|u|)+\eta_{R}^{\prime}(|u|)|u|\right)w\text{d}x.$
And then
$\displaystyle\left<\Psi^{\prime}_{\epsilon,R}\left(u_{n}\right)-\Psi_{\epsilon,R}^{\prime}\left(u_{0}\right),u_{n}-u_{0}\right$>
(T1) $\displaystyle=\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}(\nabla u_{n}\nabla(u_{n}-u_{0})+\tilde{V}(x)u_{n}(u_{n}-u_{0}))\text{d}x\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{n}|)u_{n})\text{d}x$
(T2) $\displaystyle\quad~-\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}(\nabla u_{0}\nabla(u_{n}-u_{0})+\tilde{V}(x)u_{0}(u_{n}-u_{0}))\text{d}x\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{0}|)u_{0})\text{d}x$
(T3) $\displaystyle\quad~+\frac{1}{\epsilon}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}f(\epsilon\eta_{R}(|u_{n}|)u_{n})\left(\eta_{R}(|u_{n}|)+\eta_{R}^{\prime}(|u_{n}|)|u_{n}|\right)(u_{n}-u_{0})\text{d}x$
(T4) $\displaystyle\quad~-\frac{1}{\epsilon}\eta_{R}(\|u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}f(\epsilon\eta_{R}(|u_{0}|)u_{0})\left(\eta_{R}(|u_{0}|)+\eta_{R}^{\prime}(|u_{0}|)|u_{0}|\right)(u_{n}-u_{0})\text{d}x.$
We now proceed to provide term-by-term estimates for $(\text{T}1)$ through $(\text{T}4)$. We begin by estimating $(\text{T}1)$, for which we obtain the following estimate:
$\left|(\text{T}1)\right|$ $=\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}(\nabla u_{n}\nabla(u_{n}-u_{0})+\tilde{V}(x)u_{n}(u_{n}-u_{0}))\text{d}x\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{n}|)u_{n})\text{d}x$
$=\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}|\nabla(u_{n}-u_{0})|^{2}+\tilde{V}(x)(u_{n}-u_{0})^{2})\text{d}x\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{n}|)u_{n})\text{d}x$
$\quad+\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}(\nabla u_{0}\nabla(u_{n}-u_{0})+\tilde{V}(x)u_{0}(u_{n}-u_{0}))\text{d}x\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{n}|)u_{n})\text{d}x$
$\leq\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\|u_{n}-u_{0}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{n}|)u_{n})\text{d}x$
$\quad+o_{n}(1)\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}F(\epsilon\eta_{R}(|u_{n}|)u_{n})\text{d}x$
$\leq\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\|u_{n}-u_{0}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\int_{\mathbb{R}^{3}}\left(\delta\epsilon^{2}\eta_{R}^{2}\left(\left|u_{n}\right|\right)u_{n}^{2}+C_{\delta,R}\epsilon^{p}\eta_{R}^{p}\left(\left|u_{n}\right|\right)\left|u_{n}\right|^{p}\right)\text{d}x$
$\quad+o_{n}(1)\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}\left(\delta\epsilon^{2}\eta_{R}^{2}\left(\left|u_{n}\right|\right)u_{n}^{2}+C_{\delta,R}\epsilon^{p}\eta_{R}^{p}\left(\left|u_{n}\right|\right)\left|u_{n}\right|^{p}\right)\text{d}x$
$\leq\delta\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\|u_{n}-u_{0}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x$
$\quad+C_{\delta,R}\epsilon^{p-2}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\|u_{n}-u_{0}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\int_{\mathbb{R}^{3}}\left|u_{n}\right|^{p}\text{d}x$
$\quad+(\delta\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x+C_{\delta,R}\epsilon^{p-2}\eta_{R}^{\prime}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}\left|u_{n}\right|^{p}\text{d}x)o_{n}(1)$
$\leq\delta R^{2}\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}+C_{\delta,R}\epsilon^{p-2}\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}+(\delta R^{2}+C_{\delta,R}\epsilon^{p-2})o_{n}(1),$
where we use the fact that $u_{n}\to u_{0}$ weakly in $H^{1}_{V}(\mathbb{R}^{3})$. Take $\epsilon(R)=\min\{\epsilon^{\prime}(R),\frac{\delta_{0}}{R},\left(\frac{1}{4C_{\delta,R}}\right)^{1/(p-2)}\}$ with $\delta=\frac{1}{4R^{2}}$, for any $\epsilon\in(0,\epsilon(R))$ we can deduce that
$|(\text{T}1)|\leq\frac{1}{2}\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}+o_{n}(1).$
Next, by using $u_{n}\to u_{0}~\text{weakly~in}~H_{V}^{1}\left(\mathbb{R}^{3}\right)$ and the similar method in Lemma 5, we obtain that
$|(\text{T}2)|=o_{n}\left(1\right).$
We now proceed to $(\text{T}3)$ and $(\text{T}4)$. Without loss of generality, we assume that $\epsilon(R)\in(0,1)$, for any $\epsilon\in(0,\epsilon(R))$, one has
$\left|(\text{T}3)\right|$ $\leq\frac{C_{R}}{\epsilon}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\left|\int_{\mathbb{R}^{3}}f(\epsilon\eta_{R}(|u_{n}|)u_{n})(u_{n}-u_{0})\text{d}x\right|$
$\leq C_{R}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}\left[\delta\eta_{R}\left(\left|u_{n}\right|\right)|u_{n}|+C_{\delta,R}\epsilon^{p-2}\eta_{R}^{p-1}\left(\left|u_{n}\right|\right)|u_{n}|^{p-1}\right]|u_{n}-u_{0}|\text{d}x$
$\leq\delta C_{R}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}|u_{n}||u_{n}-u_{0}|\text{d}x$
$~~~~+C_{\delta,R}\epsilon^{p-2}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\int_{\mathbb{R}^{3}}|u_{n}|^{p-1}|u_{n}-u_{0}|\text{d}x.$
$\leq\delta C_{R}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\|u_{n}\|_{L^{2}(\mathbb{R}^{3})}\|u_{n}-u_{0}\|_{L^{2}(\mathbb{R}^{3})}$
$~~~~+C_{\delta,R}\epsilon^{p-2}\eta_{R}(\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})\left(\int_{\mathbb{R}^{3}}|u_{n}|^{p}\text{d}x\right)^{\frac{p-1}{p}}\left(\int_{\mathbb{R}^{3}}\left|u_{n}-u_{0}\right|^{p}\text{d}x\right)^{\frac{1}{p}}$
$\leq\delta C_{R}\|u_{n}-u_{0}\|_{L^{2}(\mathbb{R}^{3})}+C_{\delta,R}\|u_{n}-u_{0}\|_{L^{p}(\mathbb{R}^{3})}.$
Then it follows from $u_{n}\to u_{0}$ strongly in $L^{q}(\mathbb{R}^{3})$ with $q\in[2,6)$ that $|(\text{T}3)|=o_{n}\left(1\right).$ Finally, it is easy to see that $|(\text{T}4)|=o_{n}\left(1\right).$
Combining these results, we have proved the above claim, and then one has
$\|u_{n}-u_{0}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}=o_{n}\left(1\right),$
which implies $u_{n}\to u_{0}$ strongly in $H_{V}^{1}\left(\mathbb{R}^{3}\right)$. ∎